Two cars get closer by 8m every second while travelling in the opposite directions. They get closer by 0.8m while travelling in the same direction.what are the speeds of two cars?

Grade:9

2 Answers

Shailendra Kumar Sharma
6 years ago
Assume A and B be the cars Now we can use concept of relative velocityWhen they are heading each otherVAB = VA- VB Distance traveled in 1,secAnd (VA+ VB)*1= 8And same directionVA-VB= .8SoVA= 4.4VB =3.6
M
6 years ago
Va-Vb=0.8 Va+Vb=8.°.Now we get 2 eqn solve them and Va=4.4m/s and Vb=3.6m/s hope that this is a helpful ans...................,....